OpenStudy (anonymous):

A number pattern has 75 as its first term. The rule for the pattern is subtract 6. What is the sixth term? a.)39 , b.)45 . c.) 51 , d.)69

OpenStudy (texaschic101):

an = a1 + (n - 1) d n is the term you are looking for, so its 6 a1 is the first number, so it is 75 d is the common difference, so it is -6 now we sub a6 = 75 + (6 - 1) * -6 a6 = 75 + 5 * -6 a6 = 75 + (-30) a6 = 75 - 30 a6 = 45 6th term is 45 You might want to look at this... http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

thank you texaschic101 for your help

OpenStudy (texaschic101):

no problem :)